# Tunneling dynamics of ${}^{164}$y supersolids and droplets

### v1

@@ III Initial states in the tilted double-well
Increasing $\epsilon_{{\rm dd}}$ while maintaining the confinement and dipolar direction eventually renders the system self-bound, characterized by a negative chemical potential. For instance, when $\epsilon_{{\rm dd}}\approx 1.4$ for the elongated single-well (and similarly, for $\epsilon_{{\rm dd}}=1.49$ in the case of the circular single-well) we observe the formation of a periodic, density-modulated pattern in the left well [^1] Concerning a state of the dBEC that is entirely trapped in one of the wells, the transition from the SF towards the SS and DL phases can also be seen in the momentum distribution where in contrast to the single-peak structure of the SF additional higher-lying ones accumulate in both the SS and droplet regimes.. The individual density peaks222 The number of the generated density humps increases for larger atom number while keeping all other system parameters fixed as it was also discussed e.g. in Ref. Poli et al. (2021). These structures appear in the transverse direction as elongated filaments. are phase-coherently inter-linked by a coherent background SF and a SS state forms. For larger $\epsilon_{{\rm dd}}$, the dipolar interactions become dominant. This leads to a vanishing SF background and a simultaneous stronger spatial localization of the aforementioned density humps, see Fig. 1(c), (f). In this droplet phase, the self-bound character is related to exhibiting large negative energy or equivalently chemical potential [Fig. 2(a)]. Figure 1(c) depicts a characteristic DL density distribution in an elongated geometry. However, in the planar geometry, instead, a central droplet is surrounded by a ring-shaped SF configuration, see Fig.1(f), as discussed in Bisset et al. (2016).
@@
Figure 4: Time-evolution of the population imbalance, $\Delta\mathcal{N}(t)$, between the left and right wells for (a) different initial tilts $D$ (see legend) at fixed $\epsilon_{{\rm dd}}=1.394$ referring to a SS and (b) distinct relative interactions, $\epsilon_{{\rm dd}}$ (see legend) when $D=7\hbar\omega_{x}/l_{{\rm osc}}$. It is evident that a decreasing tilt strength slows down tunneling, whilst for a constant $D$ population imbalance at long evolution times is suppressed only within the SF phase and becomes maximal for a droplet array. (c) $\Delta\mathcal{N}(t)$ in the SS case with $\epsilon_{{\rm dd}}=1.456$ and using an initial $D=7\hbar\omega_{x}/l_{{\rm osc}}$ for several barrier heights $V_{D}$ (see legend). An increasing $V_{D}$ leads to a macroscopic self-trapping regime. The dBEC consists of $N=40000$ ${}^{164}$y atoms confined in a quasi-1D double-well. The tilt strength $D$ of the double-well is measured in units of $\hbar\omega_{x}/l_{{\rm osc}}$.
@@ IV Dynamics in an elongated double-well
To explore the dynamical properties of the dBEC in a double-well, we first consider an elongated geometry along the x-direction. The system is quenched by instantaneously (at $t=0$) switching-off the external tilt, from a value $D>0$ to $D=0$. Naturally, the strength of the tilt potential is related to the initial velocity imparted on the dBEC directly after performing the $D=0$ quench. Namely, a larger initial $D$ refers to an increasing velocity. This can be readily verified by determining the velocity of the center-of-mass of the dBEC, $v(t=0)\equiv v_{0}$, provided in Fig. 2(b) for various interactions as a function of $D$. The velocity, herein, is defined as the time-derivative of the center-of-mass coordinate $X_{{\rm CM}}(t)=\int_{-x_{0}}^{x_{0}}\mathrm{dx}~{}xn_{{\rm 1D}}(x,t)$, where $n_{{\rm 1D}}(x,t)=\int\mathrm{dydz}~{}n(x,y,z,t)$ denotes the one-dimensional (1D) integrated density and ±$\pm x_{0}/2$ is the location of the employed hard-wall boundaries along the $x$-direction. These are, of course, chosen sufficiently wide such that they do not affect our results. The dependence of $v_{0}$ on the interaction strength is clearly seen. For larger $\epsilon_{{\rm dd}}$ values and a fixed $D$, i.e. towards the droplet regime, the velocity decreases, reflecting the rigidity of the state.
@@
As a starting point, we leverage a relatively large tilt strength, $D\geq 10\hbar\omega_{x}/l_{{\rm osc}}$, rendering the impact of the initial velocity $v_{0}$ (which is enhanced here, Fig. 2(b)) irrelevant. This enables us to unravel the ensuing interaction effects on the dynamics. Ramping-down $D$ leads to coherent oscillations of the dBEC with a frequency that is independent of the relative interactions $\epsilon_{{\rm dd}}$, see Fig. 3(a1)-(a3) and Fig. 4(a). Notice, however, that the oscillation amplitude is reduced for increasing $\epsilon_{{\rm dd}}$, see also the discussion below and Fig. 4(b). In the case of a SF ($\epsilon_{{\rm dd}}\approx 1.31$) the entire dBEC coherently moves between the left and right wells with constant amplitude, and period $T_{{\rm Rabi}}\approx 55\rm ms$, in analogy to what is known for regular BECs Smerzi et al. (1997); Albiez et al. (2005). This response will be dubbed macroscopic collective tunneling regime.
A similar dynamical behavior is also observed for a SS [Fig. 3(a2) for $\epsilon_{{\rm dd}}=1.397$] and a droplet array [Fig. 3(a3) for $\epsilon_{{\rm dd}}=1.49$]. Initially being localized in the left well, after quenching, the SS or DL crystal structures move towards the right side, and during this process come into close proximity with each other, particularly in the vicinity of the potential barrier (see Figs. 3 (a2) and (a3). Subsequently, the crystals move into the right well, and as the crystals reach the right well’s edge, they separate from each other again. This process repeats periodically. The maximum separation between the single crystals at the well’s edge is more pronounced for larger $\epsilon_{{\rm dd}}$ due to the dominant dipolar interactions, which enhance repulsion among the density humps in the transverse plane [see Figs. 3(a3) and (b3)].
@@
A reduction of the tilt strength (or equivalently the initial velocity, Fig. 2(b)) leads to significant alterations of the dynamical behavior. Comparing Fig. 3(a1)-(a3) with Fig. (b1)-(b3) reveals a different response, displaying intricate tunneling dynamics. At short evolution times, $t\approx\tau_{D_{{\rm in}}=7}$, the dBEC oscillates back and forth between the wells. However, owing to the reduced velocity at $t=0$, the dBEC gradually starts to be partially transmitted to the other well while a relatively small fraction is reflected back. As more particles are reflected in the course of the evolution, individual density humps become less prominent both in the SS [Fig. 3(b2)] and the DL [Fig. 3(b3)]. Eventually, after four oscillation periods ($t>>\tau_{D_{{\rm in}=7}}$) a certain population remains trapped within each well and tunneling locks. Indeed, employing $D=7\hbar\omega_{x}/l_{{\rm osc}}$ which still ensures the confinement of the dBEC within the left well at $t=0$, we observe that the smaller initial velocity (as compared to the case with $D=10\hbar\omega_{x}/l_{{\rm osc}}$) results in a gradual suppression of the inter-well tunneling, see also $\Delta\mathcal{N}(t)$ in Fig. 4(a). A further decrease of $D$ favors an appreciable part of the population to initially also reside in the right well. This eventually leads to suppression of the tunneling due to the relatively weak exerted momentum by the quench, see the values of $\Delta\mathcal{N}(t)$ for $D=2\hbar\omega_{x}/l_{{\rm osc}}$ for example in Fig. 4(a). This resembles a type of self-trapping, as it has been also realized in regular BECs Milburn et al. (1997); Albiez et al. (2005); Levy et al. (2007); Gati and Oberthaler (2007).
For larger $\epsilon_{\rm dd}$, different droplet lattice configurations can be realized in the left well compared to the right well. In this case, the individual crystals or SFs undergo multifrequency intrawell oscillations of varying amplitude induced by the tunneling of the background SF, see also the explicit time-evolution in vid (2023a). We note here that removing the small offset between the wells is essentially equivalent to imparting a small kick velocity to the crystal structures. The vibrational patterns emanate from the activation of the different underlying normal modes, as demonstrated in Mukherjee and Reimann (2023).
@@
V Dynamics in a double-well with circular single traps
Let us now investigate how the above-described phenomena depend on the dimensionality of the system, and turn to the planar geometry with circular single wells. The dBEC is initialized in the tilted double-well assembling in different configurations according to the value of $\epsilon_{{\rm dd}}$ as depicted in Fig. 1(d)-(f). As before, the time-evolution is induced via quenching the tilt strength $D$ from a finite value down to $D=0$. Similar to the elongated geometry [Fig. 2(b)], the initial velocity, $v_{0}$, exhibits a linear increase for larger $D$ and fixed $\epsilon_{{\rm dd}}$, while being reduced for larger $\epsilon_{{\rm dd}}$ and constant $D$ (not shown). The overall dynamical behavior is characterized through the two-dimensional (2D) integrated density defined as $n_{{\rm 2D}}(x,y,t)=\int\mathrm{dz}~{}n(x,y,z,t)$ and the respective population imbalance $\Delta\mathcal{N}(t)=(1/N)[\int_{-x_{0}/2}^{0}\int_{-y_{0}}^{y_{0}}\mathrm{dxdy}~{}n_{{\rm{2D}}}(x,y,t)-\int_{0}^{x_{0}/2}\int_{-y_{0}}^{y_{0}}\mathrm{dxdy}~{}n_{{\rm{2D}}}(x,y,t)]$, with ±$\pm x_{0}/2$ [±$\pm y_{0}/2$] denoting the position of the hard-walls in the $x$ [$y$] direction.
@@ V Dynamics in a double-well with circular single traps
Considering a large energy offset at $t=0$, quantified by $D=10\hbar\omega_{x}/l_{{\rm osc}}$, we find that irrespective of $\epsilon_{{\rm dd}}$ the entire dBEC undergoes regular oscillations between the left and right wells with constant amplitude and frequency throughout the time evolution. This behavior can be directly seen in the densities of a SF [SS] depicted in Fig. 6(a1)-(a6) [Fig. 7(a1)-(a6)] but also in the corresponding inter-well population imbalance provided in Fig. 8(a) exemplary for the SF case. Focusing on the evolution of the SF state with $\epsilon_{{\rm dd}}=1.31$, there are two main distinct dynamical stages: At early evolution times, the SF coherently oscillates back and forth without any noticeable density deformation. While later on [Fig. 6(a5), (a6)], the density distortion gradually enhances, likely due to the trigger of underlying collective modes Ronen et al. (2006); Guo et al. (2019); Natale et al. (2019); Tanzi et al. (2019b) after multiple collisions of the condensate with the central barrier.
@@
Figure 8: Dynamics of the inter-well population imbalance, $\Delta\mathcal{N}(t)$, in the quasi-2D double-well when considering (a) various tilt strengths $D$ (see legend) and $\epsilon_{{\rm dd}}=1.31$ corresponding to a SF or (b) using different $\epsilon_{{\rm dd}}$ (see legend) and $D=7\hbar\omega_{x}/l_{{\rm osc}}$. A reduced energy offset results to a gradual suppression of tunneling. Moreover, $\Delta\mathcal{N}(t)$ in the region of $D$ where tunneling locks vanishes within the SF phase but it remains finite in the SS. The evolution of the quasi-2D dBEC with $N=40000$ ${}^{164}$y atoms is triggered by quenching the initial potential offset $D$ to zero.
@@ VI Conclusions &\&& perspectives
We investigated the ground state and the tunneling dynamics of a dBEC trapped in an initially tilted double-well potential encompassing both elongated and circular trap geometries. To model these systems and their dynamical behavior, we employed the extended Gross-Pitaevskii framework that incorporates first-order quantum corrections. The different phases are initialized by tuning the relative strength ratio between dipolar and short-range interactions while maintaining the tilt strength. The inter-well energy offset imposed by the tilt strength effectively modifies the trapping volume of the dBEC. Thus, even for fixed interactions, when changing the tilt strengths a transition from a SF to SS and DL state may occur, decreasing the chemical potential to negative values. The tilt value, furthermore, governs the initial center-of-mass velocity of the dBEC. For larger offsets and fixed interactions, the velocity monotonically increases. Conversely, for a specific tilt and varying interactions, the velocity decreases as the system transitions from the SF phase towards the SS and DL phases, evincing the rigidity of the latter.
@@
There is a variety of possible extensions of our results in future endeavors. An imperative prospect is to testify the validity of the eGPE to adequately describe all the emergent tunneling channels. Along these lines, it is worth to unravel the participation of possible interband tunneling processes especially in the few- to many-body crossover relying, for instance, on ab-initio methods Cao et al. (2017); Haldar and Alon (2019). Additionally, the inclusion of temperature and its impact on the discussed tunneling properties is worth pursuing. A first step towards this direction could be to utilize a stochastic Gross-Pitaevskii model as it was done in Bland et al. (2022a). Exploiting dynamical frustration events, e.g., by ramping a lattice potential in SS and DL Halperin et al. (2023), to design exotic tunneling processes is another intriguing direction. The spontaneous generation of nonlinear wave structures stemming from the counterflow dynamics of two initially separated dBEC clouds is certainly desirable. Furthermore, a detailed study of the excitation of corresponding surface modes in dBECs (for example using parametric driving as it was achieved in regular BECs Engels et al. (2007); Kwon et al. (2021)) constitutes another interesting future prospect.

### v2

@@ III Initial states in the tilted double-well
Increasing $\epsilon_{{\rm dd}}$ while maintaining the confinement and dipolar direction eventually renders the system self-bound, characterized by a negative chemical potential. For instance, when $\epsilon_{{\rm dd}}\approx 1.4$ for the elongated single-well (and similarly, for $\epsilon_{{\rm dd}}=1.49$ in the case of the pancake-like double-well) we observe the formation of a periodic, density-modulated pattern in the left well [^1] Concerning a state of the dBEC that is entirely trapped in one of the wells, the transition from the SF towards the SS and DL phases can also be seen in the momentum distribution where in contrast to the single-peak structure of the SF additional higher-lying ones accumulate in both the SS and droplet regimes.. The individual density peaks222 The number of the generated density humps increases for larger atom number while keeping all other system parameters fixed as it was also discussed e.g. in Ref. Poli et al. (2021). These structures appear in the transverse direction as elongated filaments. are phase-coherently inter-linked by a coherent background SF and a SS state forms. For larger $\epsilon_{{\rm dd}}$, the dipolar interactions become dominant. This leads to a vanishing SF background and a simultaneous stronger spatial localization of the aforementioned density humps, see Fig. 1(c), (f). In this droplet phase, the self-bound character is related to exhibiting large negative energy or equivalently chemical potential [Fig. 2(a)]. Figure 1(c) depicts a characteristic DL density distribution in an elongated geometry. However, in the planar geometry, instead, a central droplet is surrounded by a ring-shaped SF configuration, see Fig.1(f), as discussed in Bisset et al. (2016).
@@
Figure 4: Time-evolution of the population imbalance, $\Delta\mathcal{N}(t)$, between the left and right wells for (a) different initial tilts $D$ (see legend) at fixed $\epsilon_{{\rm dd}}=1.394$ referring to a SS and (b) distinct relative interactions, $\epsilon_{{\rm dd}}$ (see legend) when $D=7\hbar\omega_{x}/l_{{\rm osc}}$. It is evident that a decreasing tilt strength slows down tunneling, whilst for a constant $D$ population imbalance at long evolution times is suppressed only within the SF phase and becomes maximal for a droplet array. (c) $\Delta\mathcal{N}(t)$ in the SS case with $\epsilon_{{\rm dd}}=1.456$ and using an initial $D=7\hbar\omega_{x}/l_{{\rm osc}}$ for several barrier heights $V_{D}$ (see legend). An increasing $V_{D}$ leads to a macroscopic self-trapping regime. The dBEC consists of $N=40000$ 164Dy atoms confined in a quasi-1D double-well. The tilt strength $D$ of the double-well is measured in units of $\hbar\omega_{x}/l_{{\rm osc}}$.
@@ IV Dynamics in an elongated double-well
To explore the dynamical properties of the dBEC in a double-well, we first consider an elongated geometry along the x-direction. The system is quenched by instantaneously (at $t=0$) switching-off the external tilt, from a value $D>0$ to $D=0$. For the impact of a time-dependent ramp-down of the tilt on the tunneling dynamics and the persistence of the discussed tunneling regimes for different ramp-rates see Appendix B. Naturally, the strength of the tilt potential is related to the initial velocity imparted on the dBEC directly after performing the $D=0$ quench. Namely, a larger initial $D$ refers to an increasing velocity. This can be readily verified by determining the velocity of the center-of-mass of the dBEC, $v(t=0)\equiv v_{0}$, provided in Fig. 2(b) for various interactions as a function of $D$. The velocity, herein, is defined as the time-derivative of the center-of-mass coordinate $X_{{\rm CM}}(t)=\int_{-x_{0}}^{x_{0}}\mathrm{dx}~{}xn_{{\rm 1D}}(x,t)$, where $n_{{\rm 1D}}(x,t)=\int\mathrm{dydz}~{}n(x,y,z,t)$ denotes the one-dimensional (1D) integrated density and ±$\pm x_{0}/2$ is the location of the employed hard-wall boundaries along the $x$-direction. These are, of course, chosen sufficiently wide such that they do not affect our results. The dependence of $v_{0}$ on the interaction strength is clearly seen. For larger $\epsilon_{{\rm dd}}$ values and a fixed $D$, i.e. towards the droplet regime, the velocity decreases, reflecting the rigidity of the state.
@@
As a starting point, we leverage a relatively large tilt strength, $D\geq 10\hbar\omega_{x}/l_{{\rm osc}}$, rendering the impact of the initial velocity $v_{0}$ (which is enhanced here, Fig. 2(b)) irrelevant. Indeed, for $D\geq 10\hbar\omega_{x}/l_{{\rm osc}}$, the entire dipolar gas is localized in the left double-well rendering the quench-induced tunneling process independent of the tilt strength. This enables us to unravel the ensuing interaction effects on the dynamics. Ramping-down $D$ leads to coherent oscillations of the dBEC with a frequency that is weakly dependent on the relative interactions $\epsilon_{{\rm dd}}$, see Fig. 3(a1)-(a3) and Fig. 4(b). Notice, however, that the oscillation amplitude is reduced for increasing $\epsilon_{{\rm dd}}$, see also the discussion below and Fig. 4(b). In the case of a SF ($\epsilon_{{\rm dd}}\approx 1.31$) the entire dBEC coherently moves between the left and right wells with constant amplitude, and period [^3] Notice that $T_{{\rm Rabi}}\sim 2\pi/(\omega_{x})$ for large $D$, meaning that the tunneling is essentially a collective center-of-mass oscillation, i.e. the dipole mode. However, a more accurate estimation of the tunneling period for all $D$ examined herein would require knowledge of the ensuing energy gap among the wells. In this context, a larger gap enforced by increasing tilt strength leads to a smaller period. $T_{{\rm Rabi}}\approx 55\rm$, in analogy to what is known for regular BECs Smerzi et al. (1997); Albiez et al. (2005). This response will be dubbed macroscopic collective tunneling regime.
A similar dynamical behavior is also observed for a SS [Fig. 3(a2) for $\epsilon_{{\rm dd}}=1.397$] and a droplet array [Fig. 3(a3) for $\epsilon_{{\rm dd}}=1.49$]. Initially being localized in the left well, after quenching, the SS or DL crystal structures move towards the right side, and during this process come into close proximity with each other, particularly in the vicinity of the potential barrier (see Figs. 3 (a2) and (a3). Subsequently, the crystals move into the right well, and as they reach the right well’s edge, they separate from each other again. This process repeats periodically. The maximum separation between the single crystals at the well’s edge is more pronounced for larger $\epsilon_{{\rm dd}}$ due to the dominant dipolar interactions, which enhance repulsion among the density humps in the transverse plane [see Figs. 3(a3) and (b3)].
@@
A reduction of the tilt strength (or equivalently the initial velocity, Fig. 2(b)) leads to significant alterations of the dynamical behavior. Comparing Fig. 3(a1)-(a3) with Fig. 3(b1)-(b3) reveals a different response, displaying intricate tunneling dynamics. At short evolution times, $t\approx\tau_{D_{{\rm in}}=7}$, the dBEC oscillates back and forth between the wells. However, owing to the reduced velocity at $t=0$, the dBEC gradually starts to be partially transmitted to the other well while a relatively small fraction is reflected back. As more particles are reflected in the course of the evolution, individual density humps become less prominent both in the SS [Fig. 3(b2)] and the DL [Fig. 3(b3)]. Eventually, after four oscillation periods ($t>>\tau_{D_{{\rm in}=7}}$) a certain population remains trapped within each well and tunneling locks. Indeed, employing $D=7\hbar\omega_{x}/l_{{\rm osc}}$ which still ensures the confinement of the dBEC within the left well at $t=0$, we observe that the smaller initial velocity (as compared to the case with $D=10\hbar\omega_{x}/l_{{\rm osc}}$) results in a gradual suppression of the inter-well tunneling, see also $\Delta\mathcal{N}(t)$ in Fig. 4(a). A further decrease of $D$ favors an appreciable part of the population to initially also reside in the right well. This eventually leads to suppression of the tunneling due to the relatively weak exerted momentum by the quench, see the values of $\Delta\mathcal{N}(t)$ for $D=2\hbar\omega_{x}/l_{{\rm osc}}$ for example in Fig. 4(a). This resembles a type of self-trapping, as it has been also realized in regular BECs Milburn et al. (1997); Albiez et al. (2005); Levy et al. (2007); Gati and Oberthaler (2007).
For larger $\epsilon_{\rm dd}$, different droplet lattice configurations can be realized in the left well compared to the right well. In this case, the individual crystals or SSs undergo multifrequency intrawell oscillations of varying amplitude induced by the tunneling of the background SF, see also the explicit time-evolution in vid (2023a). We note here that removing the small offset between the wells is essentially equivalent to imparting a small kick velocity to the crystal structures. The vibrational patterns emanate from the activation of the different underlying normal modes, as demonstrated in Mukherjee and Reimann (2023). The main features of the above-described tunneling behavior remain robust (at least up to certain times) even in the presence of three-body losses, which are customarily present in experiments, as showcased in Appendix A. However, in this case the tunneling is suppressed for longer evolution times and e.g. the SS or DL character is lost due to atom losses.
@@
V Dynamics in a pancake double-well
Let us now investigate how the above-described phenomena depend on the dimensionality of the system, and turn to the pancake-like double-well. The dBEC is initialized in the tilted double-well assembling in different configurations according to the value of $\epsilon_{{\rm dd}}$ as depicted in Fig. 1(d)-(f). As before, the time-evolution is induced via quenching the tilt strength $D$ from a finite value down to $D=0$. Similar to the elongated geometry [Fig. 2(b)], the initial velocity, $v_{0}$, exhibits a linear increase for larger $D$ and fixed $\epsilon_{{\rm dd}}$, while being reduced for larger $\epsilon_{{\rm dd}}$ and constant $D$ (not shown). The overall dynamical behavior is characterized through the two-dimensional (2D) integrated density defined as $n_{{\rm 2D}}(x,y,t)=\int\mathrm{dz}~{}n(x,y,z,t)$ and the respective population imbalance $\Delta\mathcal{N}(t)=(1/N)[\int_{-x_{0}/2}^{0}\int_{-y_{0}}^{y_{0}}\mathrm{dxdy}~{}n_{{\rm{2D}}}(x,y,t)-\int_{0}^{x_{0}/2}\int_{-y_{0}}^{y_{0}}\mathrm{dxdy}~{}n_{{\rm{2D}}}(x,y,t)]$, with ±$\pm x_{0}/2$ [±$\pm y_{0}/2$] denoting the position of the hard-walls in the $x$ [$y$] direction.
@@ V Dynamics in a pancake double-well
Considering a large energy offset at $t=0$, quantified by $D=10\hbar\omega_{x}/l_{{\rm osc}}$, we find that irrespective of $\epsilon_{{\rm dd}}$ the entire dBEC undergoes regular oscillations between the left and right wells with constant amplitude and frequency throughout the time evolution. This behavior can be directly seen in the densities of a SF [SS] depicted in Fig. 6(a1)-(a6) [Fig. 7(a1)-(a6)] but also in the corresponding inter-well population imbalance provided in Fig. 8(a) exemplary for the SF case. Focusing on the evolution of the SF state with $\epsilon_{{\rm dd}}=1.31$, there are two main distinct dynamical stages: At early evolution times, the SF coherently oscillates back and forth characterized by a Rabi frequency $T_{{\rm Rabi}}\approx 24\rm ms\approx 2\pi/\omega_{x}$ and without experiencing any noticeable density deformation. While later on [Fig. 6(a5), (a6)], the density distortion gradually enhances, likely due to the trigger of underlying collective modes Ronen et al. (2006); Guo et al. (2019); Natale et al. (2019); Tanzi et al. (2019b) after multiple collisions of the condensate with the central barrier.
@@
Figure 8: Dynamics of the inter-well population imbalance, $\Delta\mathcal{N}(t)$, in the quasi-2D double-well when considering (a) various tilt strengths $D$ (see legend) and $\epsilon_{{\rm dd}}=1.31$ corresponding to a SF or (b) using different $\epsilon_{{\rm dd}}$ (see legend) and $D=7\hbar\omega_{x}/l_{{\rm osc}}$. A reduced energy offset results to a gradual suppression of tunneling. Moreover, $\Delta\mathcal{N}(t)$ in the region of $D$ where tunneling locks vanishes within the SF phase but it remains finite in the SS. The evolution of the quasi-2D dBEC with $N=40000$ 164Dy atoms is triggered by quenching the initial potential offset $D$ to zero.
@@ VI Conclusions &\&& perspectives
We investigated the ground state and the tunneling dynamics of a dBEC trapped in an initially tilted double-well potential encompassing both elongated and pancake trap geometries. To model these systems and their dynamical behavior, we employed the extended Gross-Pitaevskii framework that incorporates first-order quantum corrections. The different phases are initialized by tuning the relative strength ratio between dipolar and short-range interactions while maintaining the tilt strength. The inter-well energy offset imposed by the tilt strength effectively modifies the trapping volume of the dBEC. Thus, even for fixed interactions, when changing the tilt strengths a transition from a SF to SS and DL state may occur, decreasing the chemical potential to negative values. The tilt value, furthermore, governs the initial center-of-mass velocity of the dBEC. For larger offsets and fixed interactions, the velocity monotonically increases. Conversely, for a specific tilt and varying interactions, the velocity decreases as the system transitions from the SF phase towards the SS and DL phases, evincing the rigidity of the latter.
@@
There is a variety of possible extensions of our results in future endeavors. An imperative prospect is to testify the validity of the eGPE to adequately describe all the emergent tunneling channels or even examine geometries with a few lattice sites Wilsmann et al. (2018); Grün et al. (2022). Along these lines, it is worth to unravel the participation of possible interband tunneling processes especially in the few- to many-body crossover relying, for instance, on ab-initio methods Cao et al. (2017); Haldar and Alon (2019). In this context, it would be possible to achieve an accurate characterization of the induced tunneling pathways in terms of the relevant energy gaps and possibly reveal higher-order, e.g. inter-band, tunneling channels. Additionally, the inclusion of temperature and its impact on the discussed tunneling properties is worth pursuing. A first step towards this direction could be to utilize a stochastic Gross-Pitaevskii model as it was done in Bland et al. (2022a). Exploiting dynamical frustration events, e.g., by ramping a lattice potential in SS and DL Halperin et al. (2023), to design exotic tunneling processes is another intriguing direction. The spontaneous generation of nonlinear wave structures stemming from the counterflow dynamics of two initially separated dBEC clouds is certainly desirable. Furthermore, a detailed study of the excitation of corresponding surface modes in dBECs (for example using parametric driving as it was achieved in regular BECs Engels et al. (2007); Kwon et al. (2021)) constitutes another interesting future prospect.